OpenStudy (anonymous):

Calculate the discriminant. x2 – 2x + 4 = 0 A. –12 B. 20 C. 12 D. 0

OpenStudy (anonymous):

@jagr2713 @Luigi0210

pooja195 (pooja195):

b^2−4(a)(c) list the abc values

pooja195 (pooja195):

\[\huge x^2-2x+4=0\]

pooja195 (pooja195):

a=? b=? c=?

OpenStudy (anonymous):

okay1

OpenStudy (anonymous):

a=x^2 b=2x and c=4

pooja195 (pooja195):

close enough :) a=1 b=2 c=4

pooja195 (pooja195):

\[\huge2^2-4(1)(4)\]

OpenStudy (anonymous):

okayyy

pooja195 (pooja195):

solve

OpenStudy (anonymous):